# Separate the Wheat from the Chaff: A Post-Hoc Approach to Safety Re-Alignment for Fine-Tuned Language Models

### v2

@@ Abstract
Although large language models (LLMs) achieve effective safety alignment at the time of release, they still face various safety challenges. A key issue is that fine-tuning often compromises the safety alignment of LLMs. Previous post-hoc methods make safety modifications to the model parameters as a whole, which can result in performance degradation. We propose a method named IRR (Identify, Remove, and Recalibrate for Safety Realignment) that performs safety realignment for LLMs. The core of IRR is to identify and remove unsafe delta parameters from the fine-tuned models, while recalibrating the retained parameters. We evaluate the effectiveness of IRR across various datasets, including both full fine-tuning and LoRA methods. Our results demonstrate that IRR significantly enhances the safety performance of fine-tuned models, while maintaining their performance on downstream tasks. The source code is available at: https://github.com/pikepokenew/IRR.
@@ 4 Experimental Setup
We conducted experiments using both full fine-tuning and LoRA (Hu et al., 2022). The results for full fine-tuning are presented in the main text, while LoRA results are in Appendix B. Experiments for more scenarios can be found in Appendix F and G.
@@ Appendix E Details about Llama-3-8B-Instruct experiment
Appendix F IRR with Different Safety Vectors
In the main text, the IRR method uses the same safety vector as Safe LoRA and RESTA. To explore the effectiveness of the IRR method with different safety vectors, we sampled an additional 1,000 harmful question-answer pairs from Beavertails (Ji et al., 2024) to train a new safety vector using LoRA. We present the results of the IRR and RESTA methods using a new safety vector on the model fine-tuned on GSM8K (Cobbe et al., 2021) with LoRA.
As shown in Figure 13, the IRR maintains better safety and downstream task performance compared to the RESTA when using the new safety vector. This indicates that the choice of safety vector does not affect the effectiveness of the IRR.
Appendix G IRR against Malicious Fine-tuning
To further investigate the effectiveness of IRR, we examined the scenario of mixing malicious instruction data into the fine-tuning dataset. We sampled 1,000 harmful question-answer pairs from Beavertails (Ji et al., 2024) and mixed them with the GSM8K (Cobbe et al., 2021) dataset for fine-tuning, keeping other settings the same.
We have added a special configuration to the IRR method called IRRmore. This configuration allows for the removal of other delta parameters in ascending order of their safety importance scores, after all delta parameters that cause safety interference have been removed.
Figure 14: Results of IRR against malicious fine-tuning.
As shown in Figure 14, in the case of malicious fine-tuning, the IRR method still effectively enhances safety while maintaining performance on downstream tasks compared to the RESTA method. Additionally, the IRR method can be combined with data processing methods (Zhao et al., 2023a; Yang et al., 2024) before fine-tuning and parameter update constraints (Huang et al., 2024c; Qi et al., 2024) applied during fine-tuning to achieve better safety improvements.
Appendix H Computational Complexity of IRR
@@
Appendix I Time Consumption of IRR
@@ I.1 Time for the IRR Method
I.2 Estimating Time Efficiency for Scaling to Larger Models
@@
Appendix J Limited GPU Memory Resources

### v3

@@ Abstract
Although large language models (LLMs) achieve effective safety alignment at the time of release, they still face various safety challenges. A key issue is that fine-tuning often compromises the safety alignment of LLMs. To address this issue, we propose a method named IRR (Identify, Remove, and Recalibrate for Safety Realignment) that performs safety realignment for LLMs. The core of IRR is to identify and remove unsafe delta parameters from the fine-tuned models, while recalibrating the retained parameters. We evaluate the effectiveness of IRR across various datasets, including both full fine-tuning and LoRA methods. Our results demonstrate that IRR significantly enhances the safety performance of fine-tuned models on safety benchmarks, such as harmful queries and jailbreak attacks, while maintaining their performance on downstream tasks. The source code is available at: https://github.com/pikepokenew/IRR.
@@ 4 Experimental Setup
We conducted experiments using both full fine-tuning and LoRA (Hu et al., 2022). The results for full fine-tuning are presented in the main text, while LoRA results are in the Appendix B.
@@ 5.4 Efficacy of IRR Across Models
5.5 IRR with New Safety Vector
To study the effectiveness of IRR with different safety vectors, we sampled an additional 1,000 harmful question-answer pairs from Beavertails (Ji et al., 2024) and trained a new safety vector using LoRA. We present the results of IRR and RESTA with the new safety vector on models fine-tuned with LoRA on GSM8K (Cobbe et al., 2021).
As shown in Figure 8, IRR maintains better safety and downstream task performance than RESTA, even with the new safety vector. This indicates that the choice of safety vector does not affect IRR’s effectiveness.
Figure 8: Results of IRR using the new safety vector.
5.6 IRR against Harmful Fine-tuning
To further study IRR’s effectiveness, we conducted harmful fine-tuning by mixing harmful data. We sampled 1,000 harmful question-answer pairs from Beavertails (Ji et al., 2024) and combined them with the GSM8K dataset for LoRA fine-tuning, keeping other settings unchanged. We also introduced a special IRR configuration, IRRmore, which removes additional delta parameters in ascending order of safety importance after removing all safety interference delta parameters.
The results from Figure 9 show that, in the harmful fine-tuning setting, the IRR method improves safety while maintaining higher performance on downstream tasks compared to the RESTA method, demonstrating its effectiveness.
@@ 6 Conclusion
Figure 9: Results of IRR against Harmful fine-tuning.
@@ 9 Potential Risks
10 Acknowledgements
We thank the anonymous reviewers for their comments and suggestions. And we thank Weixiang Zhao, Bicheng Wang, Hao Yang for the suggestions on the writing of this work. This work was supported by the National Natural Science Foundation of China (NSFC) via grant 62441614 and 62176078.
@@ Appendix E Details about Llama-3-8B-Instruct experiment
Appendix F IRR with New Safety Vector
We report the performance of the IRR method using the new safety vector in a jailbreak scenario. The results in Figure 15 show that the IRR method achieves better safety and downstream task performance in defending against jailbreak attacks.
Appendix G Computational Complexity of IRR
@@
Appendix H Time Consumption of IRR
@@ H.1 Time for the IRR Method
H.2 Estimating Time Efficiency for Scaling to Larger Models
@@
Appendix I Limited GPU Memory Resources